OpenStudy (anonymous):

Find a number that is between 2/7 and 4/11?

OpenStudy (mertsj):

You could find the average of the two numbers. That would be between them.

OpenStudy (mertsj):

Or you could write them with a common denominator and then you could tell what numbers are between them.

OpenStudy (anonymous):

thank you i'll try that!

OpenStudy (mertsj):

yw. Good luck

OpenStudy (anonymous):

I suppose 3/14 could be between there?

OpenStudy (unklerhaukus):

i think Mertsj's method of finding a common denominator is the best method

OpenStudy (mertsj):

2/7=.2857 3/14=.2143 4/11=.3636

OpenStudy (mertsj):

Do you think 3/14 is between the 2

OpenStudy (anonymous):

hmm. 24/77= 0.3118 ??

OpenStudy (mertsj):

That seems to work, doesn't it?

OpenStudy (mertsj):

2/7=22/77 24/77 4/11=28/77

OpenStudy (anonymous):

I guess it does. I'm going to go with that one. Thanks again!

OpenStudy (mertsj):

yw

OpenStudy (unklerhaukus):

good work

OpenStudy (anonymous):

<3